.The price for a good A has risen from 175 rub. to 210 rub. The demand for a good B has increased from 5400 units to 7100 units. Calculate the cross- price elasticity of demand? 


1
Expert's answer
2021-09-24T19:03:01-0400

"E=\frac{\delta Q_B}{\delta P_A}\times\frac{P_A}{Q_B}"


"E=\frac{7100-5400}{210-175}\times \frac{175}{5400}"


"\frac{1700}{35}\times\frac{175}{5400}=1.5741"


The cross price elasticity of demand, E is 1.5741